Project

General

Profile

2018-06-12-MEMOMIPO » History » Revision 6

Revision 5 (Rogers, Chris, 12 June 2018 16:07) → Revision 6/7 (Rogers, Chris, 12 June 2018 16:11)

h1. 2018-06-12-MEMOMIPO 

 h2. Joint MICE Experiment Management Office and MICE International Project Office 

 h3. 12th June 2018; 15:30 BST 

 "Phone details":http://mice.iit.edu/phonebridge.html 

 h4. Agenda 

 # *Introduction: KL* 
 # *Notes of previous meeting and status of actions: All* 
 ** *PS:* Generate two-pager to close-out common fund for next PSAG. 
 ** *DR:* Continue to push towards a persistence plan for the MICE IIT WWW site. 
 ** *DR:* Coordinate migration of mail server from IIT to jiscmail. 
 # *Project Manager's report: CW* 
 # *Software and computing status: DR* 
 # *Paper status and plans: CR* 
 # *AOB* 

 h4. Notes 

 Present: 
 Chris Rogers 
 Peter Garbinicius 
 Colin Whyte 
 Durga Rajaram 

 # *Introduction * Introduction - KL* KL is in another meeting, CR to chair 
 

 # *Notes of previous meeting and status of actions: All* 
 ** *PS:* Generate two-pager to close-out common fund for next PSAG. 
 *** No report - stands 
 ** *DR:* Continue to push towards a persistence plan for the MICE IIT WWW site. 
 *** Purchase order has been placed for webserver at IIT. Static mirror at RAL to be coordinated with Chris Brew. 
 ** *DR:* Coordinate migration of mail server from IIT to jiscmail. 
 *** KL will migrate EB and CB mailing lists to jiscmail 
 *** CR will migrate analysis mailing list to jiscmail 
 *** DR will migrate software mailing list to jiscmail 
 *** Other mailing lists are inactive and will be archived 
 # *Project Manager's report: CW* 
 ** See slides 
 ** Since VC 
 *** Tony will go into the synchrotron at next opportunity to check a few survey things 
 *** ISIS have offered resource to help clear MICE hall 
 *** Shipping container for US equipment now expected to arrive w/c 25th June; Alan Br likely to be available in that week 
 # *Software and computing status: DR* 
 ** V.Pec has generated a set of new TOF calibrations. TOF bad channel patch has also been brough in. DR running some tests, in time for next release. 
 ** 2017/03 geometry has wrong alignment corrections; fix is in hand. 
 ** Global track matching for backtracking from TKD is now in test. Will not make next release. 
 ** Discussion regarding "cuts" routines; they should be run by default; DR asked CR to check the default values are reasonable 
 ** Handover of computing project 
 *** Archiving mostly done 
 *** Needs some one to manage any data issues and new users (for GRID) 
 *** Durga needs to hand over geometry and "all those things". Some discussion of the details. *Action: DR* to make a list. 
 # *Paper status and plans: CR* 
 ** No update since VC 
 # *AOB* 
 ** No AOB 

 h4. Summary of actions: 

 * *PS:* Generate two-pager to close-out common fund for next PSAG. 
 * *DR:* List the tasks that need to be handed over for software and computing. 
 h3. Back to MEMO: 

 [[memo:|Back]] 

 h3. Back to MIPO: 

 [[mipo:|Back]] 

 h2. Documents